# MEGA-Bench[Uncaptioned image]: Scaling Multimodal Evaluation to over 500 Real-World Tasks

### v1

@@ E Exact Text
D.4 Breakdown results on the application dimension
@@
GPT-4o (0513) (OpenAI, 2024a) 50.3 70.1 61.4 44.0 61.0 54.9 33.2 52.8
Claude-3.5-Sonnet (Anthropic, 2024) 51.9 65.9 55.1 47.5 58.1 53.0 33.8 51.3
Gemini-1.5-Pro-002 (Google, 2024b) 43.5 53.8 57.2 41.2 58.2 52.5 33.4 51.2
Gemini-1.5-Flash-002 (Google, 2024b) 40.4 46.3 51.2 33.7 60.1 48.0 25.2 45.7
GPT-4O Mini (OpenAI, 2024b) 34.6 56.4 54.0 32.9 51.8 43.5 24.2 35.5
Qwen2-VL-72B (Alibaba, 2024) 43.7 57.5 51.6 31.2 49.7 53.4 31.2 44.9
InternVL2-Llama3-76B (Chen et al., 2024d) 29.5 42.9 46.3 28.7 47.4 42.0 21.3 30.0
LLaVA-OneVision-72B (Li et al., 2024a) 23.2 30.7 43.6 31.6 48.1 38.3 18.2 31.7
Qwen2-VL-7B (Alibaba, 2024) 32.7 42.7 42.8 25.6 42.5 39.9 20.0 29.9
Pixtral 12B (Mistral, 2024) 25.7 42.9 38.1 24.2 50.2 38.9 13.6 31.3
@@
Phi-3.5-Vision (Abdin et al., 2024) 21.9 22.3 33.3 17.6 39.5 31.6 8.9 21.9
MiniCPM-V2.6 (Yao et al., 2024) 15.3 26.6 33.2 16.5 37.8 29.1 11.7 25.7
@@ I Information-Extraction
D.5 Breakdown results on the visual input number dimension
@@
GPT-4o (0513) (OpenAI, 2024a) 56.6 49.1 45.0 47.5 52.7 53.2
Claude-3.5-Sonnet (Anthropic, 2024) 53.6 49.3 44.2 46.3 52.9 50.9
Gemini-1.5-Pro-002 (Google, 2024b) 50.3 45.5 48.9 39.1 53.0 50.3
Gemini-1.5-Flash-002 (Google, 2024b) 44.3 42.0 42.3 33.7 43.0 49.0
GPT-4O Mini (OpenAI, 2024b) 46.3 37.0 24.7 33.6 42.6 45.5
Qwen2-VL-72B (Alibaba, 2024) 49.1 45.2 36.7 31.0 53.6 49.9
InternVL2-Llama3-76B (Chen et al., 2024d) 41.4 31.5 24.4 20.3 34.5 39.6
LLaVA-OneVision-72B (Li et al., 2024a) 34.8 34.2 25.0 20.7 27.9 44.5
Qwen2-VL-7B (Alibaba, 2024) 37.6 33.0 26.4 19.4 37.3 41.1
Pixtral 12B (Mistral, 2024) 37.1 31.0 25.8 19.7 16.4 41.0
InternVL2-8B (Chen et al., 2024d) 30.0 25.3 17.7 15.4 19.8 34.8
Phi-3.5-Vision (Abdin et al., 2024) 27.8 28.5 20.2 12.5 14.2 24.7
MiniCPM-V2.6 (Yao et al., 2024) 26.3 22.3 17.9 14.0 23.5 35.3
@@
Appendix E Detailed Inspection of Model Behaviours on MEGA-Bench
@@
Figure 14 to Figure 33 shows the case study for samples from different tasks. We use distinct colors to highlight the tags in each task sample. We borrow the error case analysis template from MMMU (Yue et al., 2024a) while adding the keywords information of MEGA-Bench. We mainly focus on the flagship proprietary models.
@@
Source: MVBench (Li et al., 2024d) and STAR (Wu et al., 2024)
@@
Appendix F Detailed Task Information
In Appendix F, we list data source details for every task in our benchmark. We also list the output format and metrics to help better understand each task’s form.
@@ Appendix F Detailed Task Information
Vlnqa Egocentric Navigation Video Data collected from VLN-CE (Krantz et al., 2020) and the task is adapted from MVBench (Li et al., 2024d), the question and answer are adapted by human annotator Contextual Simple String Match
@@
Web Action Grounding Data collected from VisualWebBench (Liu et al., 2024), and the questions and answers are adapted by human annotator MC Exact String Match
@@
Perception-Test Object Shuffle Video Data collected from VLN-CE (Krantz et al., 2020) and the task is adapted from MVBench (Li et al., 2024d), the question and answer are adapted into single choice by human annotator MC Simple String Match
@@
Perception Test Video Character Order Data collected from Perception Test (Patraucean et al., 2024) and the task is adapted from MVBench (Li et al., 2024d), the question and answer are adapted into single answer string by human annotator Contextual Simple String Match
@@
CLEVER Moving Direction Video Video data are collected from MVBench (Li et al., 2024d). Questions and answers are adapted for the contextual formatted output format Contextual Multi Ref Phrase
@@
ClEVRER Video Moving Object Property Recognition The videos are collected from MVBench (Li et al., 2024d), the questions and answers are adapted to test the understanding of physical property and dynamics Contextual Multi Ref Phrase
@@
ClEVR Object Existence Video Videos are collected from MVBench (Li et al., 2024d). Questions and answers are slightly adapted MC Simple String Match
@@
CLEVRER Video Moving Object Count Video data are collected from MVBench (Li et al., 2024d). Questions and answers are adapted for the direct numerical output Numerical Exact String Match
@@
Extract Webpage Headline Images were collected from VisualWebBench (Liu et al., 2024). Questions and answers were adapted by the annotator Contextual Simple String Match
@@
Sceneqa Scene Transition Video Video data are collected from MVBench (Li et al., 2024d). Questions and answers are adapted by the annotator into open-ended format Open GPT-4o as Judge
@@
Perception Test Video Action Count Video data are collected from MVBench (Li et al., 2024d). Questions and answers are adapted by the annotator into direct numerical output format Numerical Exact String Match
@@
Web Action Prediction Data were collected from VisualWebBench (Liu et al., 2024) MC Exact String Match
@@
Sta Action Localization Video Video data are collected from MVBench (Li et al., 2024d). Questions and answers are repurposed for the contextual formatted output format Contextual Simple String Match
@@
Red Teaming Racial Images and labels come from the Red Teaming benchmark (Li et al., 2024e). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
Red Teaming Captcha Images and labels come from the Red Teaming benchmark (Li et al., 2024e). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
Red Teaming Politics Images and labels come from the Red Teaming benchmark (Li et al., 2024e). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
@@
Red Teaming Visual Order B Images and labels come from the Red Teaming benchmark (Li et al., 2024e). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
Red Teaming Celebrity Images and labels come from the Red Teaming benchmark (Li et al., 2024e). Questions were created by a human annotator. or generated by GPT-4 Open GPT-4o as Judge
@@
Red Teaming Jailbreak Images and labels come from the Red Teaming benchmark (Li et al., 2024e). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
Red Teaming Visual Order A Images and labels come from the Red Teaming benchmark (Li et al., 2024e). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
@@
Red Teaming Visualmisleading Images and labels come from the Red Teaming benchmark (Li et al., 2024e). Questions were created by a human annotator. Open GPT-4o as Judge
@@
Appendix G Author Contribution Statement
@@
Yuansheng Ni helped organize the Appendix and polished §A to §E.

### v2

@@ E Exact Text
E.4 Breakdown results on the application dimension
@@
Claude-3.5-Sonnet (1022) (Anthropic, 2024b) 51.7 65.9 56.6 47.6 61.2 55.6 39.9 55.1
GPT-4o (0513) (OpenAI, 2024a) 50.3 70.6 61.4 44.0 61.0 55.1 33.2 52.8
Claude-3.5-Sonnet (0620) (Anthropic, 2024a) 51.9 66.6 55.1 47.5 58.1 53.2 33.8 51.3
Gemini-1.5-Pro-002 (Google, 2024b) 43.5 54.2 57.2 41.2 58.2 52.5 33.4 51.2
Gemini-1.5-Flash-002 (Google, 2024b) 40.4 46.6 51.2 33.7 60.1 48.0 25.2 45.7
GPT-4o mini (OpenAI, 2024b) 34.6 56.7 54.0 32.9 51.8 43.6 24.2 35.5
Qwen2-VL-72B (Alibaba, 2024) 43.7 58.1 51.7 31.2 49.7 53.6 31.2 44.9
InternVL2-Llama3-76B (Chen et al., 2024d) 29.5 43.1 46.3 28.7 47.4 42.2 21.3 30.0
LLaVA-OneVision-72B (Li et al., 2024a) 23.2 30.8 43.6 31.6 48.1 38.4 18.2 31.7
NVLM-72B (Dai et al., 2024) 23.9 22.8 37.2 24.5 18.9 30.2 8.0 24.9
Qwen2-VL-7B (Alibaba, 2024) 32.7 42.7 42.8 25.6 42.5 40.0 20.0 29.9
Pixtral 12B (Mistral, 2024) 25.7 43.0 38.1 24.2 50.2 38.9 13.6 31.3
Aria-MoE-25B (Li et al., 2024d) 28.5 38.3 41.0 26.2 39.7 37.8 14.3 29.7
@@
Phi-3.5-Vision (Abdin et al., 2024) 21.9 22.4 33.3 17.6 39.5 31.6 8.9 21.9
MiniCPM-V2.6 (Yao et al., 2024) 15.3 26.7 33.2 16.5 37.8 29.2 11.7 25.7
@@ I Information-Extraction
E.5 Breakdown results on the visual input number dimension
@@
Claude-3.5-Sonnet (1022) (Anthropic, 2024b) 56.4 48.8 48.3 46.3 59.1 49.5
GPT-4o (0513) (OpenAI, 2024a) 56.7 49.1 45.0 47.5 53.4 53.2
Claude-3.5-Sonnet (0620) (Anthropic, 2024a) 53.7 49.3 44.2 46.3 54.1 50.9
Gemini-1.5-Pro-002 (Google, 2024b) 50.3 45.5 48.9 39.1 53.7 50.3
Gemini-1.5-Flash-002 (Google, 2024b) 44.3 42.0 42.3 33.7 43.7 49.0
GPT-4o mini (OpenAI, 2024b) 46.3 37.0 24.7 33.6 43.1 45.5
Qwen2-VL-72B (Alibaba, 2024) 49.2 45.2 36.7 31.0 54.7 49.9
InternVL2-Llama3-76B (Chen et al., 2024d) 41.5 31.5 24.4 20.3 34.8 39.6
LLaVA-OneVision-72B (Li et al., 2024a) 34.8 34.2 25.0 20.7 28.1 44.5
NVLM-72B (Dai et al., 2024) 36.8 23.3 3.8 0.0 0.0 0.0
Qwen2-VL-7B (Alibaba, 2024) 37.7 33.0 26.4 19.4 37.5 41.1
Pixtral 12B (Mistral, 2024) 37.1 31.0 25.8 19.7 16.6 41.0
Aria-MoE-25B (Li et al., 2024d) 35.8 27.3 19.8 21.1 27.1 42.9
InternVL2-8B (Chen et al., 2024d) 30.1 25.3 17.7 15.4 19.9 34.8
Phi-3.5-Vision (Abdin et al., 2024) 27.8 28.5 20.2 12.5 14.3 24.7
MiniCPM-V2.6 (Yao et al., 2024) 26.3 22.3 17.9 14.0 23.6 35.3
@@
Appendix F Detailed Inspection of Model Behaviours on MEGA-Bench
@@
Figure 14 to Figure 33 shows the case study for samples from different tasks. We use distinct colors to highlight the tags in each task sample. We borrow the error case analysis template from MMMU (Yue et al., 2024a) while adding the keywords information of MEGA-Bench. We mainly focus on the flagship proprietary models. The Claude-3.5 in these figures refers to the Claude-3.5-Sonnet (0620) model.
@@
Source: MVBench (Li et al., 2024e) and STAR (Wu et al., 2024)
@@
Appendix G Detailed Task Information
In Appendix G, we list data source details for every task in our benchmark. We also list the output format and metrics to help better understand each task’s form.
@@ Appendix G Detailed Task Information
Vlnqa Egocentric Navigation Video Data collected from VLN-CE (Krantz et al., 2020) and the task is adapted from MVBench (Li et al., 2024e), the question and answer are adapted by human annotator Contextual Simple String Match
@@
Web Action Grounding Data collected from VisualWebBench (Liu et al., 2024a), and the questions and answers are adapted by human annotator MC Exact String Match
@@
Perception-Test Object Shuffle Video Data collected from VLN-CE (Krantz et al., 2020) and the task is adapted from MVBench (Li et al., 2024e), the question and answer are adapted into single choice by human annotator MC Simple String Match
@@
Perception Test Video Character Order Data collected from Perception Test (Patraucean et al., 2024) and the task is adapted from MVBench (Li et al., 2024e), the question and answer are adapted into single answer string by human annotator Contextual Simple String Match
@@
CLEVER Moving Direction Video Video data are collected from MVBench (Li et al., 2024e). Questions and answers are adapted for the contextual formatted output format Contextual Multi Ref Phrase
@@
ClEVRER Video Moving Object Property Recognition The videos are collected from MVBench (Li et al., 2024e), the questions and answers are adapted to test the understanding of physical property and dynamics Contextual Multi Ref Phrase
@@
ClEVR Object Existence Video Videos are collected from MVBench (Li et al., 2024e). Questions and answers are slightly adapted MC Simple String Match
@@
CLEVRER Video Moving Object Count Video data are collected from MVBench (Li et al., 2024e). Questions and answers are adapted for the direct numerical output Numerical Exact String Match
@@
Extract Webpage Headline Images were collected from VisualWebBench (Liu et al., 2024a). Questions and answers were adapted by the annotator Contextual Simple String Match
@@
Sceneqa Scene Transition Video Video data are collected from MVBench (Li et al., 2024e). Questions and answers are adapted by the annotator into open-ended format Open GPT-4o as Judge
@@
Perception Test Video Action Count Video data are collected from MVBench (Li et al., 2024e). Questions and answers are adapted by the annotator into direct numerical output format Numerical Exact String Match
@@
Web Action Prediction Data were collected from VisualWebBench (Liu et al., 2024a) MC Exact String Match
@@
Sta Action Localization Video Video data are collected from MVBench (Li et al., 2024e). Questions and answers are repurposed for the contextual formatted output format Contextual Simple String Match
@@
Red Teaming Racial Images and labels come from the Red Teaming benchmark (Li et al., 2024f). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
Red Teaming Captcha Images and labels come from the Red Teaming benchmark (Li et al., 2024f). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
Red Teaming Politics Images and labels come from the Red Teaming benchmark (Li et al., 2024f). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
@@
Red Teaming Visual Order B Images and labels come from the Red Teaming benchmark (Li et al., 2024f). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
Red Teaming Celebrity Images and labels come from the Red Teaming benchmark (Li et al., 2024f). Questions were created by a human annotator. or generated by GPT-4 Open GPT-4o as Judge
@@
Red Teaming Jailbreak Images and labels come from the Red Teaming benchmark (Li et al., 2024f). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
Red Teaming Visual Order A Images and labels come from the Red Teaming benchmark (Li et al., 2024f). Questions were created by a human annotator or generated by GPT-4. Open GPT-4o as Judge
@@
Red Teaming Visualmisleading Images and labels come from the Red Teaming benchmark (Li et al., 2024f). Questions were created by a human annotator. Open GPT-4o as Judge
@@
Appendix H Author Contribution Statement
@@
Yuansheng Ni helped organize the Appendix and polished §B to §F.